OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 - 16x + 57

OpenStudy (yolo_king):

What do you think?

OpenStudy (anonymous):

I have no idea

OpenStudy (triciaal):

lost everything try again HINT think about completing the square general form ax^2 + bx + c vertex form a(x-h)^2 + k h = -b/2a and k = -(-c + h^2)

OpenStudy (yolo_king):

If you were to follow ^^^ You would get (8,-7)